OpenStudy (help!!!!):

Determine r so that the line that passes through the given points has the given slope

OpenStudy (help!!!!):

(r,5) (-2,r) m=-2/11

OpenStudy (help!!!!):

like I always get the answer r=6.5 repeating but I know its wrong

OpenStudy (help!!!!):

so basically what i did was r-5\[\frac{ r-5 }{ -2-r }=\frac{ -2 }{ 11 }\]

jimthompson5910 (jim_thompson5910):

maybe you need to provide the answer as a fraction

OpenStudy (help!!!!):

like so? 6.5-5=1.5 -2-6.5=-8.5 \[\frac{ 1.5 }{ -8.5 }\]

jimthompson5910 (jim_thompson5910):

I would cross multiply 11(r-5) = -2(-2-r) then solve for r

OpenStudy (help!!!!):

thats what i've tried but i get the answer 6.5 repeating. I tried in terms of fraction but i got -3/17

jimthompson5910 (jim_thompson5910):

11(r-5) = -2(-2-r) 11r-55 = 4+2r 11r-2r = 4+55 r = ??

OpenStudy (help!!!!):

6.55555555

jimthompson5910 (jim_thompson5910):

answer as a fraction though, before you got 6.555

OpenStudy (help!!!!):

59/9?

jimthompson5910 (jim_thompson5910):

yes, r = 59/9

OpenStudy (help!!!!):

wait but if i add in 59/9 for "a" for the points and find slope, it doesnt equal to -2/11

jimthompson5910 (jim_thompson5910):

you should have (59/9 - 5)/(-2 - 59/9) = -2/11 I'm not sure what you mean by "wait but if i add in 59/9 for "a" for the points and find slope, it doesnt equal to -2/11"